### v2

# arXiv:2505.02449v2[cond-mat.mes-hall] 01 Mar 2026

Abstract
The intense simulation efforts on hole-spin qubits in germanium have so far focused primarily on singly occupied quantum dots. Here, we theoretically investigate three-hole qubits in germanium and demonstrate that their performance can rival that of single-hole qubits in both strained and unstrained systems. In particular, we find that — in the widely used quasi-circular geometry — a three-hole qubit encoding can yield enhancements of the Rabi frequencies of up to two orders of magnitude and a large advantage also in terms of quality factors.
Figure 1: (a) Pictorial representation of a single-hole state (left) and a three-hole state (right) in a QD elongated along the $x$ direction. (b) Schematic view of a MOS device, where the Ge channel has a width $L$ along the $z$ direction. The bias applied to the metal top gate determines the in-plane confining potential $V_{\parallel}$ and the linear (in $z$) contribution in the out-of-plane confining potential $V_{\perp}$.
Figure 2: In-plane particle distribution $N_{z}(x,y)$ for the ground state $|\Phi_{0}\rangle$ of the three-hole system, obtained by integrating the particle density $n({\bf r})$ along the $z$ direction. The four panels display: (a) the function $N_{z}(x,y)$; (b,c) the heavy-hole contributions; (d) the sum of the two light-hole contributions. All the plotted functions are normalized to 1, the isolines correspond to values that are multiples of 0.1. The values of the parameters are: $\hbar\omega_{x}=5\,$meV, $\hbar\omega_{y}=6\,$meV, B=0.05B=0.05\,T and $\theta=0$ (parallel field orientation).
Introduction.
Spin qubits in semiconductor quantum dots (QDs) are promising competitors for the implementation of quantum computers [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. In particular, the low natural abundance of magnetic nuclei, combined with the possibility of isotopic purification, grants long coherence times for spins in silicon (Si) and germanium (Ge) [3, 2, 16, 17, 18, 19, 20, 21, 22, 23]. In these systems, high fidelities have been achieved in all fundamental computational tasks, from one- and two-qubit gates [24, 25, 26, 27, 28, 29] to initialization and readout [30, 31, 32]. Besides, the advanced fabrication techniques and integration circuitry provide excellent scalability prospects and allow a high degree of control even at relatively high temperatures [33, 15, 34, 35, 11, 14, 36, 18].
In contrast to electrons, holes in Ge have no valley degeneracy and are characterized by strong spin-orbit coupling (SOC) which allows for all-electric manipulation [37, 38, 21, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. In addition, due to their $p$-orbital character, they are less susceptible than electrons to the hyperfine interaction [49, 6, 50, 51]. Finally, the characteristic properties of the hole-spin qubit, including the $g$ tensor and the Rabi frequencies, are electrically tunable [35, 52]. This results from the susceptibility to the electrostatic environment of the mixing between heavy-hole (HH), light-hole (LH), and split-off (SO) bands [53, 54, 55, 56, 39, 57, 12, 8, 58].
The vast majority of the literature on hole-spin qubits focuses on single-particle encodings. Multiple holes, localized in singly-occupied QDs, have also been considered for alternative qubit encodings [10, 59, 60, 61, 62, 63, 20, 64]. On the other hand, an odd number of holes in a single QD enables qubit encodings and manipulation schemes analogous to those implemented in the single-particle case [65, 66, 67]. Such a multi-hole single-dot qubit has a twofold interest. First, from an experimental point of view, it relaxes the stringent requirement of achieving the single-occupation regime. Second, and most importantly, it might result in larger and more local couplings to the control gates and, ultimately, in improved figures of merit [68]. However, despite their past and recent occurrence in experiments [11, 12, 44, 45, 69, 70], few-particle effects in semiconductor hole-spin qubits remain relatively unexplored at the theoretical level.
In this Letter, we provide a theoretical study of multi-hole qubits in Ge QDs, implemented in unstrained MOS-like devices and in Ge//Si1-xGex heterostructures (Fig. 1). Focusing on the case of three holes, we show that the main figures of merit are always at least comparable to those obtained for single holes, and significantly improved in specific regimes. More specifically, the Rabi frequency of the three-hole qubit (THQ) is shown to exceed that of a single-hole qubit (SHQ) by up to two orders of magnitude in quasi-circular QDs. This occurs without a significant modification of the $g$ factor or loss of the qubit coherence under charge noise. As the comparison of the fully interacting THQ with its noninteracting counterpart shows, this gain is caused, for the largest part, by the occupation of the excited orbitals dictated by the Pauli principle, the hole-hole interaction giving a quantitatively smaller (albeit far from negligible) contribution in that regime. Strain changes the absolute values of the manipulation frequencies, but does not substantially affect the relative performances of the single- and multi-hole qubits.
Quantum-dot model.
Single-hole states are derived through the numerical diagonalization of the 6-band $\bm{k}\cdot\bm{p}$ Hamiltonian resulting from the Luttinger-Kohn (LK) envelope-function approach [71, 72, 53, 73]. The 6 bands needed to describe single-hole states in Ge close to the top of the valence band form a $J=3/2$ quadruplet and a $J=1/2$ doublet, with $J$ being the total atomic angular momentum associated with the Bloch states at the $\Gamma$ point. These two multiplets are split at $\Gamma$ by spin-orbit coupling (SOC), which induces a gap $\Delta_{\rm SO}=290$ meV. The $J=3/2$ quadruplet, degenerate at $\Gamma$, is split into the HH ($|M|=3/2$) and the LH ($|M|=1/2$) doublets at finite $k$. The band structure around the $\Gamma$ point is determined by the LK parameters, which, for Ge, are: $\gamma_{1}=13.38$, $\gamma_{2}=4.24$ and $\gamma_{3}=5.69$. The reference frame, specified by the unit vectors $\bm{u}_{x}$, $\bm{u}_{y}$, and $\bm{u}_{z}$, is defined with respect to the crystallographic axes as follows: $\bm{u}_{x}\parallel[110]$, $\bm{u}_{y}\parallel[\bar{1}10]$, $\bm{u}_{z}\parallel[001]$.
QD1 $g^{(1)}_{\theta}$ $g^{(3)}_{\theta}$ $f^{(3)}_{R,x}$ $f^{(3)}_{R,y}$
$\theta=0^{\circ}$ 8.67 7.99 (6.25) 46.8 (239) 9.24 (27.3)
$\theta=90^{\circ}$ 0.121 0.0848 (0.104) 7.29 (3.93) 2.07 (0.0775)
QD2 $g^{(1)}_{\theta}$ $g^{(3)}_{\theta}$ $f^{(3)}_{R,x}$ $f^{(3)}_{R,y}$
$\theta=0^{\circ}$ 8.00 4.88 (4.34) 394 (322) 313 (175)
$\theta=90^{\circ}$ 0.114 0.220 (0.321) 27.1 (23.9) 11.3 (2.04)
Table 1: Values of the $g$ factors and of the Rabi frequencies (expressed in MHz) for the unstrained single- and three-hole qubits, for $\hbar\omega_{y}=6\,$meV, $B=0.05$ T, and $\left|\delta\bm{E}_{\rm R}\right|=1$ mV//nm. The value of $\hbar\omega_{x}$ is 3 and 5.55.5\,meV for QD1 and QD2, respectively. In parentheses we report the Rabi frequencies of the noninteracting THQ.
The confining potential that defines the dot model varies smoothly with respect to the unit-cell scale, and is specifically given by
$V_{\rm conf}(\bm{r})=\frac{1}{2}\frac{m_{0}}{\gamma_{1}}\left(\omega_{x}^{2}x^{2}+\omega_{y}^{2}y^{2}\right)+V_{\perp}(z)\,.$ (1)
Here, $m_{0}$ is the free-electron mass, while $\hbar\omega_{y}$ (set to 6 meV) and $\hbar\omega_{x}$ (from 3 to 5.5 meV) are the characteristic energy scales associated with the anisotropic harmonic confinement in the $xy$ plane. The out-of-plane confinement defines a quantum well, resulting from the band offset between Ge and the barrier material:
$V_{\perp}(z)\!=\!W\!+\!\frac{W}{2}\sum_{\xi=\pm 1}\tanh\left[\xi\frac{z}{s}\!-\!\frac{(1+\xi)}{2}\frac{L}{s}\right]\!+\!e_{\perp}z\,.$ (2)
Here, L=10L=10\,nm is the well width, $W$ is the barrier height, and s=0.02s=0.02\,nm a smoothing factor. The dc electric field $e_{\perp}=1\,$meV//nm, possibly associated to a top gate voltage, is attractive for holes.
The coupling to the magnetic field is introduced into the single-hole Hamiltonian through the Zeeman-Bloch term and the Peierls substitution [52, 35], with the electromagnetic vector potential expressed in the Coulomb gauge. The magnetic field is given in polar form, $\bm{B}\equiv B\,(\sin\theta\,\cos\phi,\,\sin\theta\,\sin\phi,\,\cos\theta)$, with B=0.05B=0.05\,T, $\phi=45\degree$, and a variable polar angle $\theta$.
The few-particle (three-hole) eigenstates are computed by diagonalizing the interacting Hamiltonian, within the full configuration interaction approach [74, 75]. The three-hole Slater determinants are built from a set of 64 single-particle states, which ensures the convergence of the relevant quantities. Further details on the method are provided in the Supplemental Material (SM) [76].
Single- and three-hole spin qubits.
The magnetic field breaks the Kramers degeneracy within the ground doublet, and induces the energy splitting $\Delta^{(k)}_{10}$ between the states $\left|0^{(k)}\right\rangle$ and $\left|1^{(k)}\right\rangle$, which define the SHQ ($k=1$) and the THQ ($k=3)$. This splitting can equivalently be expressed in terms of the $g$ factor $g^{(k)}_{\theta}=\Delta^{(k)}_{10}/\mu_{\rm B}B$, which depends on the orientation of $\bm{B}$ with respect to the $z$ axis, defined by $\theta$.
Figure 3: Rabi frequencies $f^{(n)}_{{\rm R},\alpha}$ of the $n$-hole qubit as functions of $\hbar\omega_{x}$, with $n=1$ (red curves) and $n=3$ (blue symbols), when the oscillating electric field (of amplitude $\left|\delta\bm{E}_{\rm R}\right|=1$ mV//nm) is oriented along the $x$ (a, c) or $y$ direction (b, d). The magnetic field amplitude is $|\bm{B}|=0.05$ T; its direction is given by $\theta=0\degree$ (a, b) or $\theta=90\degree$ (c, d), with $\phi=45\degree$.
The qubit manipulation is performed by applying an oscillating electric field $\delta\bm{E}(t)=\delta\bm{E}_{\rm R}\cos(\omega t)$, in resonance with the Larmor frequency ($\omega=2\pi f_{\rm L}$). The resulting Rabi frequency is given by:
$f_{\rm R}^{(k)}=\frac{e}{h}\Big|\delta\bm{E}_{\rm R}\cdot\langle 0^{(k)}|\hat{\bm{r}}|1^{(k)}\rangle\Big|\,$, (3)
where $\hat{\bm{r}}$ is the hole position operator, and $|\delta\bm{E}_{\rm R}|=1\,$mV/nm is the amplitude of the spatially homogeneous oscillating electric field.
As a first means to characterize the three-hole ground state $|0^{(3)}\rangle$, we compute the particle density
$n(\bm{r})=\sum_{\beta}n_{\beta}(\bm{r})=\sum_{\beta}\langle 0^{(3)}|\hat{\psi}_{\beta}^{\dagger}(\bm{r})\,\hat{\psi}_{\beta}(\bm{r})|0^{(3)}\rangle\,$, (4)
where $\beta$ runs over the six hole bands and the field operator $\hat{\psi}_{\beta}(\bm{r})$ annihilates a hole belonging to the band $\beta$ at the position $\bm{r}$. Being the particle density along the vertical direction weakly correlated with that in the planar directions [76], we focus on the in-plane particle distribution, represented by $N_{z}(x,y)\equiv\int dz\,n(\bm{r})$ (Fig. 2). The three holes clearly display a linear arrangement along the weak-confinement direction ($x$), characterized by the presence of three peaks [panel (a)], which become increasingly resolved for decreasing values of $\omega_{x}$ [76]. Besides, the minority [$M=+3/2$, panel(b)] and majority [M=3/2M=-3/2, panel(c)] HH components of $N_{z}(x,y)\equiv\int dz\,n(\bm{r})$ are clearly related to the central and external peaks, respectively: the ground state of the THQ is thus characterized by an “antiferromagnetic" ordering in the pseudospin defined by the HH component ($M=\pm 3/2$). The in-plane distribution of the LH components [panel (d)], whose weight represent less than 1% of the total, display markedly different features.
[…]
QD1 $f^{(1)}_{R,x}$ $f^{(1)}_{R,y}$ $\tau^{(1)}$ $f^{(3)}_{R,x}$ $f^{(3)}_{R,y}$ $\tau^{(3)}$
$\theta=0^{\circ}$ 57.5 27.0 5.78 23.0 5.30 5.92
$\theta=90^{\circ}$ 0.276 0.0804 1480 0.0941 0.157 1130
QD2 $f^{(1)}_{R,x}$ $f^{(1)}_{R,y}$ $\tau^{(1)}$ $f^{(3)}_{R,x}$ $f^{(3)}_{R,y}$ $\tau^{(3)}$
$\theta=0^{\circ}$ 4.48 10.7 5.93 63.7 33.9 4.50
$\theta=90^{\circ}$ 0.0223 0.0494 752 0.301 0.349 3690
[…]
While the $g$ factors of the THQ and of the SHQ show a very similar dependence on the orientation of $\bm{B}$, the Rabi frequencies differ remarkably in the two cases (Fig. 3). In fact, for all the considered orientations of $\bm{B}$ and of the oscillating electric field $\delta\bm{E}_{\rm R}$, the Rabi frequency of the THQ displays a non-monotonic dependence on the dot aspect ratio ($\omega_{x}$). Moreover, the $f_{R,\alpha}^{(3)}$ are characterized by the presence of pronounced maxima, and by regions corresponding to a large enhancement with respect to the SHQ, mostly — but not exclusively — concentrated in the quasi-circular dot regime ($\hbar\omega_{x}\gtrsim 5\,$meV). Sharp peaks in the Rabi frequencies of the THQ show up at $\hbar\omega_{x}\approx 3.8\,$meV. These features correspond to, and are most likely related to, the presence of a narrow anticrossing between the first and second excited doublet [76].
The question arises as to whether these striking differences between the THQ and the SHQ are mainly due to Coulomb interaction or to the occupation of excited single-particle states, imposed by the Pauli principle [77]. In order to discriminate between these two cases, we compare the above values of the Rabi frequencies with those obtained for the noninteracting three-hole system ($\bar{f}_{R,\alpha}$). The difference between the two cases is always remarkable and shows that the values of the Rabi frequencies are strongly affected by the Coulomb interactions. Whether the Rabi frequencies of the THQ are larger or smaller that those of the noninteracting three-hole system depends on the orientation of the magnetic field and on the degree of elongation of the QD (Table 1).
Figure 4: Dephasing time scale $\tau$ as a function of $\hbar\omega_{x}$ (with $\hbar\omega_{y}=6\,$meV), for a magnetic field of intensity B=0.05B=0.05\,T oriented (a) parallel ($\theta=0^{\circ}$) or (b) perpendicular ($\theta=90^{\circ}$) to the $z$ axis. The charge-noise induced electric field is oriented along the $z$ direction, with $\left|\bm{E}_{\rm cn}\right|=10^{-2}$ mV/nm.
Qubit dephasing.
Random fluctuations in the energy gap $\Delta_{10}$, induced by charge noise, are generally regarded as the main source of decoherence for hole-spin qubits [2]. The dephasing time scale $\tau$ can thus be defined as:
$\frac{1}{\tau^{(k)}}\equiv\frac{e}{h}\Big|\bm{E}_{\rm cn}\cdot\left(\left\langle 1^{(k)}|\hat{\bm{r}}|1^{(k)}\right\rangle-\left\langle 0^{(k)}|\hat{\bm{r}}|0^{(k)}\right\rangle\right)\Big|\,.$ (5)
Here, $\bm{E}_{\rm cn}$ is a characteristic value of the charge-noise induced electric field vector. If the field is spatially homogeneous throughout the QD region then, due to the symmetries of the confinement potential, the dominant contribution in $1/\tau$ is the one related to the $z$ component of $\bm{E}_{\rm cn}$. As in the case of the Rabi frequencies, attention should be paid to the trends and to the comparison between single- and three-hole spin qubits, rather than to the absolute values of the dephasing time scale, which depend on the assumed value of the charge-noise induced electric field.
The differences between the values of $\tau$ obtained in the single- and in the three-hole qubits are remarkable. Similarly to the case of the Rabi frequencies, the monotonic dependence of $\tau$ on the dot aspect ratio ($\omega_{x}$) that is obtained for the SHQ is replaced — in the case of the THQ — by a non-monotonic dependence, which varies qualitatively with the magnetic-field orientation (Fig. 4). In particular, we note that the dephasing time scale of the THQ is lower than that of the SHQ for a quasi-circular dot. However, the enhancement that the THQ allows in terms of Rabi frequency overcompensates such an effect, resulting in a higher value of a quality factor $Q_{x}\equiv f_{{\rm R},x}\tau$ (see below). As for the Rabi frequencies, the region corresponding to $\hbar\omega_{x}\approx 3.8\,$meV is characterized by strong variations of $\tau$, which we attribute to the effect on the ground state doublet of the narrow anticrossing affecting the two lowest excited doublets [76].
Effect of strain.
We hereafter investigate the effect of a homogeneous biaxial strain, such as the one that can result from the lattice mismatch in Ge/Si1-xGex heterostructures. The Hamiltonian term related to a small and homogeneous strain is derived, within the $\bm{k}\cdot\bm{p}$ formalism, by the Bir-Pikus theory [78, 79, 80]. In particular, we consider a diagonal uniaxial strain tensor, $\varepsilon_{\beta}^{\alpha}=\delta_{\beta}^{\alpha}\varepsilon_{\alpha}^{\alpha}$, with $\varepsilon_{x}^{x}=\varepsilon_{y}^{y}=\varepsilon_{\parallel}$ and $\varepsilon_{z}^{z}=-2\varepsilon_{\parallel}C_{12}/C_{11}$, (where $C_{11}=126$ GPa and $C_{12}=44$ GPa are the elastic constants of Ge [81]), and a typical Ge concentration $x=0.8$ in the alloy, for which the measured strain parameter is $\varepsilon_{\parallel}=-0.0063$ [82].
The presence of strain significantly affects the values of the quantities considered so far. The strain increases the energy difference between the HH and LH components, enhancing the HH character of the single-hole ground state. This results in an increased $g$-factor anisotropy, both in the SHQ and in the THQ. As to the Rabi frequencies, their values are generally reduced by the presence of strain, to a similar extent for the SHQs and the THQs. However, the presence of biaxial strain also results in a systematic increase of $\tau^{(1)}$ and $\tau^{(3)}$, as shown in Table 2 for the cases of the elongated and quasi-circular dots. Overall, the QD quality factor presents non-monotonic dependencies on the in-plane aspect ratio, with no general advantage or disadvantage resulting from the presence of strain (Fig. 5). On the other hand, the value of $Q_{x}$ obtained for the THQ largely exceeds that of the SHQ (figure inset) in the quasi-circular geometry, both for the strained and for the unstrained dot. This shows that the advantage of the three-hole encoding in terms of Rabi frequency ovecompensates the advantage of the SHQ at the level of dephasing time scale.
Figure 5: Quality factor $Q_{x}$ for the THQ as a function of $\hbar\omega_{x}$ (with $\hbar\omega_{y}=6\,$meV), for a magnetic field intensity B=0.05B=0.05\,T oriented out fo plane (parallel to the $z$ axis). Yellow and purple symbols correspond to the strained and to the unstrained dots, respectively. Inset: $Q_{x}$ vs $\hbar\omega_{x}$ for the SHQ, with the same units and the same color code as in the main plot.
Conclusions.
The various scenarios considered in this work consistently point to the fact that three-hole spin qubits in Ge perform better than their single-hole counterparts when the QD is close to the circular regime. In both cases, the largest Rabi frequencies associated to in-plane oscillating electric fields are obtained for magnetic fields oriented along the growth direction. Biaxial strain leads to an overall reduction of the Rabi frequencies, which is however balanced by a comparable increase in the characteristic dephasing time, resulting in no significant difference between strained and unstrained dots in terms of quality factor.
Acknowledgements.
We are particularly grateful to Y.-M. Niquet for an insightful comment, and we thank the members of the Niquet, Veldhorst, Rimbach-Russ, and Scappucci groups for helpful discussions. We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support (IsCb4-GERONIMO - HP10CG99PB). This work was funded by the Italian PNRR Project PE0000023-NQSTI and, as a part of NCCR SPIN, a National Centre of Competence in Research, by the Swiss National Science Foundation (grant number 225153). SB was also supported by the EU through the H2024 QLSI2 project and the Army Research Office under Award Number: W911NF-23-1-0110. DL acknowledges the Deanship of Research at KFUPM and the Quantum Center for the support received under Grant no. CUP25102 and no. INQC2500, respectively. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.